Given an=an−1−6an−2 where a0=1 a2=5

a.) list the first 10 terms of the sequence

b.) find a closed form(solve the recurrence relations) 


1
Expert's answer
2022-03-07T23:03:02-0500

Solution


Given that


"{a_n} = {a_{n - 1}} - 6{a_{n - 2}}\"


And, "{a_0} = 1" , "{a_2} =5"


For "n=2"


"{a_2} = {a_{2 - 1}} - 6{a_{2 - 2}}\\
{a_2} = {a_{1}} - 6{a_{0}}\"


using, "{a_0} = 1" , "{a_2} =5"


"5 = {a_{1}} - 6(1)\rightarrow {a_{1}}=11"


Similarly, for For "n=3"

"{a_3} = {a_{2}} - 6{a_{1}}\\
{a_3} = 5 - 6(11)\\
{a_3} =-61"


Similarly, for For "n=4"

"{a_4} = {a_{3}} - 6{a_{2}}\\
{a_4} = -61-6(5)\\
{a_4} =-91"


Continuing in the same way, we get the next few terms, which are tabulated as shown below.